OpenStudy (anonymous):

A 68 mile trip uses 2 gallons of gas. How many gallons would you use on a 110 mile trip? Round to the nearest tenth of a gallon. A. 1.2 B. 3.2 C. 3.8 D. 4.1

OpenStudy (anonymous):

Well you can cancel one of the answers right away. It has to be more than 2 gallons.

OpenStudy (anonymous):

68 mile = 2 gallons 1 mile = 2/68 gallon 110 mile = 110 * 2/ 68 gallon = 3.2 gallon ans = b 3.2 gallon

OpenStudy (anonymous):

Try not to give answers.

OpenStudy (anonymous):

ok i won't @romeo thanks for reminding :D

OpenStudy (anonymous):

You did work explaining but you should always let the asker do some of the work. Like the the last step. I think the asker knows how to divide.